### v1

@@ 5 Experiments
Here, we are interested in whether models have learned detailed knowledge about a concept. We pose the evaluation as a ranking problem with one image on one side and $n$ different texts on the other side. Besides the correct text (containing the correct keyword), we generate all possible $n-1$ negative examples using the same procedure as in § 3.1 and rank the texts with the model. We evaluate the ranking using the top-1 accuracy. This approach is conducted on the This was evaluated on the respective sub-datasets in which keywords occur. See the Supplementary for some examples.
@@
Figure 4: Fine-grained Concept Understanding vs. General Image Retrieval: Results for four different concepts trained on corresponding dataset subsets. Checkpoints are evaluated after every 10% of the data (circles); checkpoints at epoch ends are marked with the respective numbers.
@@ Results.
Results are shown in Figure 4 and exact results per epoch are displayed in the Table 3. The performance of the original OpenAI CLIP is shown with a black dot and constantly reaches the worst results. Continued pretraining on COCO massively increases the general retrieval performance for all concepts, showing the successful domain adaptation regarding this dataset. It also improves the concept understanding for all concepts except for location. We are especially interested in the trade-off between general retrieval and concept understanding for the different types of further pretrained models. The following values are based on model checkpoints at the end of the epoch.
For objects, we observe a performance increase regarding the fine-grained comprehension from 0.56 for the original OpenAI CLIP models to 0.76 when further pretraining on COCO. Using the hard negatives approach, the performance further increases by 7 to 10 percentage points depending on the hard negative proportion and training duration. Here, the general retrieval performance only drops by 1 to 2 percentage points in relation to the classical approach.
We observe a similar pattern can be seen for the concept color. Understanding of the concept improved by 11 to 15 percentage points compared to the original training process. On the other hand, the general comprehension pattern loses 3 to 4 percentage points with HN1 and slightly more with hard negative values. In this case, one hard example seems to be completely sufficient to learn concepts.
@@
We observe a similar pattern for the size concept. Further pretraining on COCO improves concept understanding by 5 to 8 percentage points to around 75%. If, on the other hand, the new learning concept is used, the result is 90%, which corresponds with an additional improvement of 15 percentage points. Depending on training time, there is no or only a 3 percentage points loss in the general image retrieval capabilities. As with location, an almost continuous improvement can be observed.
@@
Accuracy T2I R@5
Concept Epoch Orig. Clas. HN1 HN2 HN3 Orig. Clas. HN1 HN2 HN3
object 1 .56 .76 .83 .84 .86 .53 .68 .67 .67 .66
2 .56 .76 .84 .85 .85 .53 .68 .67 .67 .67
3 .56 .76 .83 .85 .86 .53 .69 .68 .68 .68
color 1 .48 .69 .81 .82 .83 .53 .64 .61 .60 .59
2 .48 .71 .82 .83 .84 .53 .64 .61 .61 .59
3 .48 .69 .82 .83 .84 .53 .64 .60 .59 .57
location 1 .58 .59 .89 .53 .65 .64
2 .58 .61 .90 .53 .65 .64
3 .58 .62 .91 .53 .65 .63
size 1 .69 .74 .90 .53 .64 .64
2 .69 .75 .90 .53 .65 .63
3 .69 .77 .90 .53 .65 .62
Object Color Size
Epoch Orig. Clas. HN1 HN2 HN3 Orig. Clas. HN1 HN2 HN3 Orig. Clas. HN1
1 .78 .85 .88 .88 .88 .62 .83 .89 .89 .89 .26 .27 .55
2 .78 .84 .86 .88 .88 .62 .83 .90 .90 .91 .26 .28 .57
3 .78 .87 .88 .89 .90 .62 .83 .90 .91 .92 .26 .30 .60
@@ 5.2 Challenge Set Results
We consider an image pair from the InpaintCOCO dataset to be correctly classified if the correct images would be more likely retrieved based on the original caption and the newly created caption. This leads to the formula,
@@
The results show that with all three concepts (object, color, size), continued pretraining improves the concept understanding. The improvements between the original OpenAI CLIP model and the further pretrained model are 6 to 9 percentage points for the object and 21 for the color concept. The improvements are less distinct for size, with a 1 to 4 percentage point gain, which aligns with the textual comprehension results displayed in Figure 3(d).
For the object concept, hard negative training brought a 7 to 10 percentage point improvement for the textual viewpoint (see Table 3). For the evaluation from the visual perspective, improvements are 2 to 4 percentage points. This relatively smaller improvement is likely because each object could be replaced with the 79 other object names during training. Still, in this evaluation, a replacement was only executed within the COCO super-category (5 to 10 object names per super-category). The differences in performance between the three models using hard negative training is $\leq 2$ percentage points.
@@
For the size concept, we see a big improvement for both perspectives when using hard negative models. From the visual perspective, there is an improvement of over 28 percentage points, and from the textual perspective (Table 3) 13 to 16 percentage points.
@@
Dataset Concept Size Orig. Clas. HN1 HN2 HN3
@@ 5.3 Evaluations on other Datasets
We further investigate the performance of our model using more VL datasets.
@@
First, we evaluate the models using general VL datasets. The investigated concepts occur with different frequencies, and for high-quality results, it is important that these concepts are understood to increase the overall performance. Therefore, we additionally investigate fine-grained concept understanding on the Flickr30k [25] and SBU Captioned Photo [11] datasets.
@@
Fine-grained concept understanding is also important in specific domains. For example, in fashion, a correct assignment of garments and colors is important, not the mere presence of colors in the image. For this analysis, we evaluated our models on the very specific datasets Fashion200K [3], NASA Earth Instagram, [^5] https://huggingface.co/datasets/nkasmanoff/nasa_earth_instagram and Old Book Illustrations. [^6] https://huggingface.co/datasets/gigant/oldbookillustrations These datasets are very heterogeneous in their appearance.
@@
All models achieve good results except the color concept on the Fashion200k dataset and object concept on Old Book Illustrations. For the former, this is because images usually show garments with a distinct color. Yet, there is little background noise or noise from irrelevant items, which can confuse the color alignment of the model in this dataset. The latter shows old-fashioned drawings in which the objects are very dissimilar to those in the COCO dataset and object list. Our approach to learn concepts works very well for the remaining evaluations.
6 Limitations and Risks
@@
Our research introduces a novel method for training CLIP aimed at incorporating concepts and representations that are challenging to learn using current approaches using weak negative samples only. In our experiments, we only used one specific CLIP model; however, we believe there is no reason why the method should not work or work systematically differently for smaller or larger CLIP-based models.
@@
We conducted training with the well-studied COCO captioning dataset, which is standard in multimodal research. The proposed method is expected to show consistent performance also using other multimodal training datasets. Notably, evaluations on out-of-domain datasets, where the model was not trained, emphasize the robustness of our approach. One essential prerequisite for our methodology to work is the presence of keywords of interest in the training corpus and language and domain knowledge to decide how the keywords should be permuted. The keyword permutation will be more difficult in languages with more complex morphology than in English. Experiments involved using four concepts with a carefully chosen set of keywords. Depending on domain-specific tasks, other keywords might be of interest (e.g., a large list of garments for the fashion domain).
@@
7 Conclusion
We introduce a robust method for enhancing fine-grained concept understanding with minimal impact on general retrieval capabilities using hard negative sampling in contrastive learning. We demonstrate that various concepts can be learned successfully in a data-efficient way with minor adjustments to text inputs. We can already observe improvements in concept understanding after continued pretraining on only 10% of our data. Furthermore, one hard negative sample per image in a batch of 64 proves to be sufficient to incorporate the concept of interest into the model.
We comprehensively evaluate our method on several datasets, including a new challenge set, InpaintCOCO. Our method outperforms classical contrastive learning on all four investigated concepts. Existing datasets often focus on linguistic perturbations or use dissimilar image pairs, precluding a structured evaluation of permuted visual concepts in isolation. To address this gap, InpaintCOCO represents the first dataset adjusting minor image parts in a controlled setting, facilitating cross-model fine-grained understanding. This ensures that the model’s output is influenced only by the object of interest and not the rest of the scene.
When using both the InpaintCOCO dataset and other domain-specific datasets, the results show that fine-grained concept understanding also generalizes to images of different styles. The method is data-efficient for the images needed and requires only a little domain knowledge for designing the hard negative prompts. It makes the method suitable for domain adaptation of image retrieval for specific use cases without the need to train with multiple objective functions or larger models.
Acknowledgements
The authors gratefully acknowledge the compute time granted on the GPU cluster Monacum One at Bundeswehr University Munich. The creation of the dataset was supported by the Charles University grant PRIMUS/23/SCI/023.
Appendix 0.A Appendix
0.A.1 Creating Hard Negative Samples
In Table 6, we specify all used permutation keywords for the concepts object, color, location and size.
Table 7 lists text samples for the fine-grained concept understanding task.
Concept
Keywords
object
[person, bicycle, car, motorbike, aeroplane, bus, train, truck, boat, traffic light, fire hydrant, stop sign, parking meter, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, backpack, umbrella, handbag, tie, suitcase, frisbee, skis, snowboard, sports ball, kite, baseball bat, baseball glove, skateboard, surfboard, tennis racket, bottle, wine glass, cup, fork, knife, spoon, bowl, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake, chair, sofa, potted plant, bed, dining table, toilet, tv monitor, laptop, mouse, remote, keyboard, cell phone, microwave, oven, toaster, sink, refrigerator, book, clock, vase, scissors, teddy bear, hair drier, toothbrush]
color
[blue, red, green, yellow, black, white, brown, gray, orange]
location
left $\leftrightarrow$ right, above $\leftrightarrow$ below, under $\leftrightarrow$ over, foreground $\leftrightarrow$ background, in front of $\leftrightarrow$ behind, back $\leftrightarrow$ front
size
large $\leftrightarrow$ small, little $\leftrightarrow$ big, tall $\leftrightarrow$ short, long $\rightarrow$ short, thin $\leftrightarrow$ fat, huge $\leftrightarrow$ tiny, giant $\rightarrow$ tiny
@@ 0.A.1 Creating Hard Negative Samples
Figure 5: Sample images from MS COCO used in Table 7.
@@
Table 7: Exemplary image descriptions that are used as text samples in the fine-grained concept understanding task for the example images displayed in Figure 5.
0.A.2 COCO training and evaluation data
@@
There are many large-scale VL datasets available to pretrain or further pretrain models (e.g., Conceptual Caption [19] or LAION-5B [18] with 3.3 million and 5 billion image-text pairs respectively). Yet, results in Figure 4 in the main paper indicate that training on a small-sized COCO dataset (even for just one epoch) is sufficient to learn the concepts of interests.
@@
a big cat is next to a small dog
a small cat is next to a big dog
the businessperson’s down fall
the businessperson’s fall down
@@
0.A.3 Challenge Set Creation
The challenge set was created by undergraduate student workers. They were provided with an interactive Python environment with which they interacted via various prompts and inputs. The description of the task and the problem of the research question was made available to them (see Figure 7). In addition to a detailed written explanation of how the tool works, they were also given “best practices,” which were created by one student and reviewed by the authors.
@@ 0.A.3 Challenge Set Creation
A subset of students had pre-existing roles within the university, while others were purposefully recruited for the designated task. The compensation for student assistants adhered to the legally stipulated wages of approx. 12 EUR $\approx$ 13 USD $\approx$ 93 CNY.

### v2

@@ 5 Experiments
Here, we are interested in whether models have learned detailed concept knowledge. We pose the evaluation as a ranking problem with one image on one side and $n$ different texts on the other side. Besides the correct text (containing the correct keyword), we generate all possible $n-1$ negative examples using the same procedure as in § 3.1 and rank the texts with the model. We evaluate the ranking using the top-1 accuracy. See Table 4 in the Appendix for some examples.
@@
Figure 4: Fine-grained Concept Understanding vs. General Image Retrieval: Results for four different concepts trained on corresponding dataset subsets. Checkpoints are evaluated after every 10% of the data (circles); checkpoints at epoch ends are marked with the respective numbers. The results are also in a table form in Table 7 in the Appendix.
@@ Results.
Results are shown in Figure 4 (and exact results per epoch are displayed in Table 7 in the Appendix). The performance of the original OpenAI CLIP is shown with a black dot and constantly reaches the worst results. Continued pretraining on COCO massively increases the general retrieval performance for all concepts, showing the successful domain adaptation regarding this dataset. It also improves the concept understanding for all concepts except for location. We are especially interested in the trade-off between general retrieval and concept understanding for the different types of further pretrained models. The following values are based on model checkpoints at the end of the epoch.
For objects, we observe a performance increase regarding the fine-grained comprehension from 0.56 for the original OpenAI CLIP models to 0.76 when further pretraining on COCO. Using the hard negatives approach, performance increases by 7 to 10 percentage points, depending on the hard negative proportion and training duration. Here, the general retrieval performance only drops by 1 to 2 percentage points in relation to the classical approach.
We observe a similar pattern can be seen for the concept color. Understanding of the concept improved by 11 to 15 percentage points compared to the original training process. On the other hand, the general comprehension pattern loses 3 to 4 percentage points with HN1 and slightly more with hard negative values. In this case, one hard example seems sufficient to learn concepts.
@@
We observe a similar pattern for the size concept. Further pretraining on COCO improves concept understanding by 5 to 8 percentage points to around 75%. On the other hand, if the new learning concept is used, the result is 90%, which corresponds to an additional improvement of 15 percentage points. Depending on training time, the general image retrieval capabilities lose no or only 3 percentage points. As with location, an almost continuous improvement can be observed.
@@
Epoch Orig. Clas. HN1 HN2 HN3
1 0.78 0.85 0.88 0.88 0.88
2 0.78 0.84 0.86 0.88 0.88
3 0.78 0.87 0.88 0.89 0.90
(a) Object
Epoch Orig. Clas. HN1 HN2 HN3
1 0.62 0.83 0.89 0.89 0.89
2 0.62 0.83 0.90 0.90 0.91
3 0.62 0.83 0.90 0.91 0.92
(b) Color
Epoch Orig. Clas. HN1
1 0.26 0.27 0.55
2 0.26 0.28 0.57
3 0.26 0.30 0.60
(c) Size
@@ 5.2 Challenge Set Results
We consider an image pair from the InpaintCOCO dataset correctly classified if the correct images would be more likely to be retrieved based on the original caption and the newly created caption. This leads to the formula,
@@
The results show that continued pretraining improves understanding of all three concepts (object, color, size). The improvements between the original OpenAI CLIP model and the further pretrained model are 6 to 9 percentage points for the object and 21 for the color concept. The improvements are less distinct for size, with a 1 to 4 percentage point gain, which aligns with the textual comprehension results displayed in Figure 4(d).
For the object concept, hard negative training brought a 7 to 10 percentage point improvement for the textual viewpoint (see Table 7). For the evaluation from the visual perspective, improvements are 2 to 4 percentage points. This relatively smaller improvement is likely because each object could be replaced with the 79 other object names during training. Still in this evaluation, a replacement was only executed within the COCO super-category (5 to 10 object names per super-category). The differences in performance between the three models using hard negative training is $\leq 2$ percentage points.
@@
For the size concept, we see a big improvement for both perspectives when using hard negative models. From the visual perspective, there is an improvement of over 28 percentage points (Table 1(c)), and from the textual perspective, 13 to 16 percentage points.
@@
Dataset Concept Count Orig. Clas. HN1 HN2 HN3
@@ 5.3 Evaluations on other Datasets
We further investigate the performance of our model using more VL datasets. First, we evaluate the models using general VL datasets. The investigated concepts occur with different frequencies, and for high-quality results, it is important that these concepts are understood to increase the overall performance. Therefore, we further investigate fine-grained concept understanding on the Flickr30k (Young et al., 2014) and SBU Captioned Photo (Ordonez et al., 2011) datasets.
@@
Fine-grained concept understanding is also important in specific domains. For example, in fashion, a correct assignment of garments and colors is important, not the mere presence of colors in the image. For this analysis, we evaluated our models on the very specific datasets Fashion200K (Han et al., 2017), NASA Earth Instagram, [^5] https://huggingface.co/datasets/nkasmanoff/nasa_earth_ instagram and Old Book Illustrations. [^6] https://huggingface.co/datasets/gigant/oldbookillustrations These datasets are very heterogeneous in their appearance.
@@
All models achieve good results except the color concept on the Fashion200k dataset and object concept on Old Book Illustrations. For the former, this is because images usually show garments with a distinct color. Yet, there is little background noise or noise from irrelevant items, which can confuse the color alignment of the model in this dataset. The latter shows old-fashioned drawings with objects very dissimilar to those in the COCO dataset. Our approach to learn concepts works very well for the remaining evaluations.
6 Conclusion
We introduce a robust method for enhancing fine-grained concept understanding with minimal impact on general retrieval capabilities using hard negative sampling in contrastive learning. We show that various concepts can be learned efficiently with minor text input adjustments. Moreover, improvements in concept understanding are observable after continued pretraining on only 10% of our data. Furthermore, one hard negative sample per image in a batch of 64 proves sufficient to incorporate the concept of interest into the model.
We comprehensively evaluate our method on several datasets, including our new challenge set. Our method outperforms classical contrastive learning on all investigated concepts. Existing datasets often focus on linguistic perturbations or use dissimilar images, precluding a structured evaluation of permuted visual concepts in isolation. To address this gap, InpaintCOCO represents the first dataset adjusting minor image parts in a controlled setting, facilitating cross-model fine-grained understanding. This ensures that the model’s output is influenced only by one object and not the rest of the scene.
The results show that fine-grained concept understanding also generalizes to images of different styles when using InpaintCOCO and domain-specific datasets. Our method is data-efficient and requires only a little domain knowledge to design the hard negatives. This makes it particularly suitable for domain adaptation in image retrieval, as well as for developing new CLIP-based models, e.g., for object detection (Minderer et al., 2023).
Acknowledgements
The authors gratefully acknowledge the compute time granted by Monacum One at Bundeswehr University Munich. The work at CUNI, including annotation payments, was supported by Charles University project PRIMUS/23/SCI/023.
Limitations and Risks
@@ Acknowledgements
Our research introduces a novel method for training CLIP aimed at incorporating concepts and representations that are challenging to learn using current approaches. In our experiments, we only used one specific CLIP model; however, we believe there is no reason why the method should not work or work systematically differently for smaller or larger CLIP models.
@@
We conducted training with the well-studied COCO captioning dataset, which is standard in multimodal research. The proposed method is expected to show consistent performance also using other multimodal training datasets. Notably, evaluations on out-of-domain datasets, where the model was not trained, emphasize the robustness of our approach. One essential prerequisite for our methodology to work is the presence of keywords of interest in the training corpus and language and domain knowledge to decide how the keywords should be replaced. The keyword substitution will be more difficult in languages with more complex morphology than in English. Experiments involved using four concepts with a carefully chosen set of keywords. Depending on domain-specific tasks, other keywords might be of interest (e.g., a large list of garments for the fashion domain).
@@
Appendix A Appendix
Creating Hard Negative Samples
In Table 3, we specify all used substitution keywords for the concepts object, color, location, and size.
Table 4 lists text samples for the fine-grained concept understanding task, which is used in § 5.1 for each concept.
Concept Keywords
object [person, bicycle, car, motorbike, aeroplane, bus, train, truck, boat, traffic light, fire hydrant, stop sign, parking meter, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, backpack, umbrella, handbag, tie, suitcase, frisbee, skis, snowboard, sports ball, kite, baseball bat, baseball glove, skateboard, surfboard, tennis racket, bottle, wine glass, cup, fork, knife, spoon, bowl, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake, chair, sofa, potted plant, bed, dining table, toilet, tv monitor, laptop, mouse, remote, keyboard, cell phone, microwave, oven, toaster, sink, refrigerator, book, clock, vase, scissors, teddy bear, hair drier, toothbrush]
color [blue, red, green, yellow, black, white, brown, gray, orange]
location left $\leftrightarrow$ right, above $\leftrightarrow$ below, under $\leftrightarrow$ over, foreground $\leftrightarrow$ background, in front of $\leftrightarrow$ behind, back $\leftrightarrow$ front
size large $\leftrightarrow$ small, little $\leftrightarrow$ big, tall $\leftrightarrow$ short, long $\rightarrow$ short, thin $\leftrightarrow$ fat, huge $\leftrightarrow$ tiny, giant $\rightarrow$ tiny
@@ Appendix A Appendix
Table 4: Exemplary image descriptions that are used as text samples in the fine-grained concept understanding task (see § 5.1) for the example images displayed in Figure 1 or Figure 3.
COCO training and evaluation data.
@@
There are many large-scale VL datasets available to pretrain or further pretrain models (e.g., Conceptual Caption (Sharma et al., 2018) or LAION-5B (Schuhmann et al., 2022) with 3.3 million and 5 billion image-text pairs respectively). Yet, results in Figure 4 indicate that training on a small-sized COCO dataset (even for just one epoch) is sufficient to learn the concepts of interest.
@@
Fine-grained VL Benchmarks.
In Table 6, we list different benchmark datasets for fine-grained understanding in VL. Some image-text samples for ARO and Winoground are displayed in Figure 5 since these datasets provide two images per sample – similar to InpaintCOCO. However, unlike InpaintCOCO, the visual representations are very different regarding the scenes presented in these datasets.
Dataset Size Perspective Samples Tasks
ARO 77k Linguistic 1 image $\leftrightarrow$ 2 texts Attributes, Relations, Order understanding
VL-CheckList 410k Linguistic 1 image $\leftrightarrow$ 2 texts Attributes, Relations, Object understanding
SVO-Probs 48k Visual 2 images $\leftrightarrow$ 1 text Relations (Verb Understanding)
InpaintCOCO 1,260 Cross-modal 2 images $\leftrightarrow$ 2 texts Attributes, Object understanding
Table 6: Datasets for fine-grained understanding in VL.
Refer to caption Refer to caption
@@
Refer to caption Refer to caption
a big cat is next to a small dog a small cat is next to a big dog
Refer to caption Refer to caption
the businessperson’s down fall the businessperson’s fall down
@@
Challenge Set Creation.
Undergraduate student workers created the challenge set. They were provided with an interactive Python environment with which they interacted via various prompts and inputs. The description of the task and the problem of the research question was made available to them (see Figure 6). In addition to a detailed written explanation of how the tool works, they were also given “best practices,” which were created by one student and reviewed by the authors.
@@
A subset of students had pre-existing roles within the university, while others were purposefully recruited for the designated task. The compensation for student assistants adhered to the legally stipulated wages in their respective countries, amounting to CZK 300.00 per hour in the Czech Republic and EUR 12.00 per hour in Germany.
@@
Experiment Results.
For numeric results from Figure 4 see Table 7. In this Table fine-grained concept understanding results (Accuracy) and COCO text-to-image retrieval results (T2I R@5) are presented. “Orig.”, “Clas.”, “HN1”, “HN2”, and “HN3” indicate the original OpenAI CLIP model, the classical further pretrained model, and models trained with 1, 2, or 3 hard negative samples.
Accuracy T2I R@5
Concept Epoch Orig. Clas. HN1 HN2 HN3 Orig. Clas. HN1 HN2 HN3
object 1 .56 .76 .83 .84 .86 .53 .68 .67 .67 .66
2 .56 .76 .84 .85 .85 .53 .68 .67 .67 .67
3 .56 .76 .83 .85 .86 .53 .69 .68 .68 .68
color 1 .48 .69 .81 .82 .83 .53 .64 .61 .60 .59
2 .48 .71 .82 .83 .84 .53 .64 .61 .61 .59
3 .48 .69 .82 .83 .84 .53 .64 .60 .59 .57
location 1 .58 .59 .89 .53 .65 .64
2 .58 .61 .90 .53 .65 .64
3 .58 .62 .91 .53 .65 .63
size 1 .69 .74 .90 .53 .64 .64
2 .69 .75 .90 .53 .65 .63
3 .69 .77 .90 .53 .65 .62
Refer to caption Refer to caption
A bird is standing on top of a car. A cat is standing on top of a car.
(a) Object
Refer to caption Refer to caption
A couple of white bathroom sinks sitting next to a toilet. A couple of black bathroom sinks sitting next to a toilet.
(b) Color
Refer to caption Refer to caption
A living room with white furniture and a small wooden table. A living room with white furniture and a huge wooden table.
(c) Size
Figure 7: InpaintCOCO samples for all concepts.